OpenStudy (anonymous):

Solve using Quadratic Formula x^2 + 7x + 4 = 0

hartnn (hartnn):

again ?

hartnn (hartnn):

what answer did u get ?

OpenStudy (anonymous):

haha i knowww. But this one has a weird square root that i can't get passed

hartnn (hartnn):

49-4*4 = 49-16 = 33

OpenStudy (anonymous):

do i leave is as the square root of 33 the whole time? or can i simplify it?

hartnn (hartnn):

nopes, can't simplify, leave it as \(\sqrt{33}\)

OpenStudy (anonymous):

Let me take another crack at it and ill let you know what i get :]

OpenStudy (anonymous):

orrr is that the end?

OpenStudy (anonymous):

Got it :P THANKS AGAIN! you'e saving a life today :]

hartnn (hartnn):

THE END is here: \(\huge\frac{-7 \pm \sqrt{33}}{2}\) oh, ok welcome :)